A partivle starts from rest with an acceleratuon if 2 metres per second square for 10 seconds.after that the particle travels for 30 seconds with constant speed and then undergoes retardation of 4 metres per second square and comes back to resy .the total distance covered by the paryicle is (1)650m(2) 700m(3)750m(4)800m

Grade:12th pass

1 Answers

Satwik modi
6 years ago
For acceleration part
u=0
a=2 m/s^2
t=10 s
 
x1=(0)t+1/2*2*100
x1=100m
 
v=0+2*10
v=20 m/s
 
For constant velocity part
x2=20*30
x2=600 m
 
For deceleration part
0=20 - 4t
t=5 s
 
x3=20*5-1/2*4*25
x3=50 m
 
Total dist.= 100+600+50=750 m